# EPW varieties as moduli spaces on ordinary GM surfaces and special GM threefolds

### v2

@@ Abstract.
We show that the double dual EPW sextic and double EPW sextic associated with a strongly smooth Gushel–Mukai surface can be realized as moduli spaces of semistable objects with respect to a stability condition on the bounded derived category of it. Also, we observe that double dual EPW surface and double EPW surface associated with a special Gushel–Mukai threefold can be realized as moduli spaces of semistable objects on its Kuznetsov component. As an application, we refine a statement of Bayer and Perry about Gushel–Mukai threefolds with equivalent Kuznetsov components for the special ones.
@@
Guehsl–Mukai varieties, moduli space, derived category
@@
Primary 18G80 14J28 14J45
The first author is a member of the INdAM group GNSAGA (2025). The second author is supported by SUSY Starting Grant No. 34000-12256019.
@@ 1. Introduction
Moduli spaces of semistable sheaves provide new methods to study the geometry of low dimension varieties and provide models to some higher dimensional varieties. Here we refer the readers to Huybrechts and Lehn [26] for a comprehensive account of this theory. On the other hand, moduli spaces of semistable sheaves on special members in a family of varieties are barely understood even for families of surfaces. Nevertheless, moduli spaces of semistable sheaves are in general only compatible with étale covers but not branched ones. These two disadvantages can be partly solved by considering instead the moduli spaces of semistable objects with respect to a notion of stability conditions introduced by Bridgeland [10].
@@ 1.1. GM varieties
An ordinary Gushel–Mukai surface is a smooth dimension transverse intersection $S=\operatorname{Gr}(2,V_{5})\cap\mathbb{P}(W)\cap Q$ where $V_{5}$ is a five-dimensional vector space, $W\subset\mathsf{\Lambda}^{2}V_{5}$ is a seven-dimensional vector subspace, and $Q\subset\mathbb{P}(W)$ is a quadratic hypersurface, such that the intersection $M_{S}:=\operatorname{Gr}(2,V_{5})\cap\mathbb{P}(W)$ is transversal. The surface $S$ is a Brill–Noether general K3 surface of degree $10$ under the ample polarization $H$ given by $\mathscr{O}_{S}(H)\mathscr{O}_{\mathbb{P}(W)}(1)|_{S}$.
The surface $S$ is called strongly smooth if the intersection $M_{S}$ is smooth. The double cover $X\rightarrow M_{S}$ branched over a strongly smooth ordinary Gushel–Mukai surface $S$ is called the special Gushel–Mukai threefold associated with $S$. Such a threefold $X$ has Picard rank $1$ and is Fano of index one and degree $10$ under the polarization $\mathscr{O}_{X}(1)$ given by pulling back $\mathscr{O}_{\mathbb{P}(W)}(1)$.
@@
of $X$, where the vector bundle $\mathscr{U}_{X}$ is induced by the tautological bundle $\mathscr{U}_{G}$ on $\operatorname{Gr}(2,V_{5})$. The associated equivariant category is equivalent to ${\textbf{{D}}}^{b}(S)$ according to [30, Section 8.2]. The dual cyclic group $\mathfrak{S}^{\vee}_{2}$ acts on ${\textbf{{D}}}^{b}(S)$ by an involutive autoequivalence $\Pi$ on ${\textbf{{D}}}^{b}(X)$ and the associated equivariant category is reversely equivalent to ${\textbf{{Ku}}}(X)$ by [19, Theorem 1.3].
According to [33, Secion 6.1], one can find a stability condition $\sigma_{S}$ on ${\textbf{{D}}}^{b}(S)$ for any strongly smooth ordinary Gushel–Mukai surface $S$ such that the moduli spaces of semistable objects with respect to $\sigma_{S}$ detect many geometry of $S$. Moreover, the stability condition $\sigma_{S}$ is compatible with the $\mathfrak{S}^{\vee}_{2}$-group action on ${\textbf{{D}}}^{b}(S)$ so that it induces a stability condition $\sigma_{X}$ on the equivariant category ${\textbf{{Ku}}}(X)$ and vice versa by virtue of [43, Section 4.3].
The forget functors between ${\textbf{{D}}}^{b}(S)$ and ${\textbf{{Ku}}}(X)$ induce covering morphisms between moduli spaces of semistable objects with respect to $\sigma_{S}$ and $\sigma_{X}$. The properties and first examples of these covering morphisms have been investigated in [33, 43]. In particular, some classic branched double covers can be recovered from these examples. This article is aim to recover some EPW varieties associated with $S$ and $X$ as moduli spaces, expanding results in [33, Section 6].
@@ 1.2. EPW varieties
The vector space $V_{6}:=H^{0}(\mathbb{P}(W),\mathscr{I}_{S}(2))$ of quadrics in $W$ containing an ordinary Gushel–Mukai surface $S$ is six-dimensional and each ordinary Gushel–Mukai surface $S$ is canonically associated with a Lagrangian subspace $A(S)\subset\mathsf{\Lambda}^{3}V_{6}$. One can construct a four-dimensional projective variety $\tilde{Y}^{\geq 1}_{A(S)}$ from $A(S)$, called the double EPW sextics associated with the Gushel–Mukai surface $S$. The annihilator $A(S)^{\perp}\subset\mathsf{\Lambda}^{3}V^{\vee}_{6}$ of the Lagrangian subspace $A(S)$ is also a Lagrangian subspace and gives another four-dimension projective variety $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ called the double dual EPW sextics associated with $S$.
The singular locus of $\tilde{Y}^{\geq 1}_{A(S)}$ is an integral normal Cohen–Macaulay surface $Y^{\geq 2}_{A(S)}$ called the EPW surface associated with $S$. The unique double cover over $Y^{\geq 2}_{A(S)}$ branched on its singular locus is called the double EPW surface and denoted by $\tilde{Y}^{\geq 2}_{A(S)}$. Similarly, one has the associated dual EPW surface $Y^{\geq 2}_{A(S)^{\perp}}$ and associated double dual EPW surface $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$.
@@ 1.3. Main results
In [33], the sextics $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ and $\tilde{Y}^{\geq 1}_{A(S)}$ are realized as moduli spaces of semistable objects with respect to $\sigma_{S}$ for a general strongly smooth ordinary Gushel–Mukai surface $S$ based on the descriptions in [37, 40]. Here we provide the modular description of $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ for any strongly smooth ordinary Gushel–Mukai surface following [18, Section 7.5].
@@
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{1})$ of semistable objects with Mukai vector $v_{1}=(1,0,-1)$ is isomorphic to the double dual EPW sextics $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ associated $S$.
However, a useful birational model for $\tilde{Y}^{\geq 1}_{A(S)}$ is not available. In this case, one can pass to a strongly smooth ordinary Gushel–Mukai surface $S^{\prime}$ which is derived equivalent to the surface $S$ under a $\mathfrak{S}_{2}^{\vee}$-equivariant equivalence and satisfies $\tilde{Y}_{A(S)}\cong\tilde{Y}_{A(S^{\prime})^{\perp}}$ using [7, 40]. Then one can deduce the following statement using the results in [33, Section 6].
@@
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{2})$ of semistable objects with Mukai vector $v_{2}=(2,-H,2)$ is isomorphic to the double EPW sextics $\tilde{Y}^{\geq 1}_{A(S)}$ associated $S$.
Consequently, the inflation functor ${\textbf{{Ku}}}(X)\rightarrow{\textbf{{D}}}^{b}(S)$ induces two double covers
@@
The classes $\kappa_{1}$ and $\kappa_{2}$ form a basis for $\textsf{{N}}({\textbf{{Ku}}}(X))$ which is a rank two free abelian group so that an equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ either preserves or exchanges them up to a sign. It is conjectured in [31] and proved in [7, Theorem 5.12] using homological projective duality that two special Gushel–Mukai threefolds $X$ and $X^{\prime}$ have equivalent Kuznetsov components if and only if one has $A(X)\cong A(X^{\prime})$ or $A(X)\cong A(X^{\prime})^{\perp}$. Using the computations in [7, 33], a refinement can be established as an application of Theorem 1.1 and Proposition 1.2.
@@
Consider two special Gushel–Mukai threefolds $X$ and $X^{\prime}$, then $A(X)\cong A(X^{\prime})$ if and only if one has ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ sending $\kappa_{1}$ to $\kappa^{\prime}_{1}$ and $\kappa_{2}$ to $\kappa^{\prime}_{2}$ up to a sign; $A(X)^{\perp}\cong A(X^{\prime})$ if and only if one has ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ sending $\kappa_{1}$ to $\kappa^{\prime}_{2}$ and $\kappa_{2}$ to $\kappa^{\prime}_{1}$ up to a sign.
@@
In the end of this article, we exploit our results together with the Brill–Noether locus identification of a special Gushel–Mukai threefold in [29] to investigate the special Gushel–Mukai threefolds sharing the same Lagrangian subspaces.
@@ A Gushel–Mukai variety $X$ is a smooth $n$-dimensional intersection
Proposition 2.4 ([20, Lemma 2.7, Lemma 2.8]).
@@
In fact, due to [8, Remark 2.19], a Gushel–Mukai surface is possibly strongly smooth only when it is ordinary in the following definition.
@@ 2.2. The Lagrangian subspaces and EPW varieties
is canonically associated with a Lagrangian subspace $A(X)\subset\mathsf{\Lambda}^{3}V_{6}$ for $V_{6}=H^{0}(\mathbb{P}(W),\mathscr{I}_{X}(2))$ according to [15, Theorem 3.10]. One can define a series of projective varieties associated with the Lagrangian subspace. These varieties are introduced and intensively studied by O’Grady.
@@
can be endowed with a scheme structure $Y_{A}^{\geq\ell}$ for any integral $\ell$ according to [37, Section 2].
@@
A non-zero element of a Lagrangian subspace $A\subset\mathsf{\Lambda}^{3}V_{6}$ is called decomposable once it can be written as the form $v_{1}\land v_{2}\land v_{3}$.
@@
    the singular locus of is the finite set $f^{-1}_{A}(Y_{A}^{\geq 3})$;
@@
    the variety is a $K3[2]$-type hyperkähler manifold once $Y_{A}^{\geq 3}=\varnothing$.
@@
    the variety $\tilde{Y}_{A}^{\geq 1}$ is simply connected;
@@
    the singular locus of is the finite set $f^{-1}_{A}(Y_{A}^{\geq 3})$;
  • (c)
    the variety is a $K3[2]$-type hyperkähler manifold once $Y_{A}^{\geq 3}=\varnothing$.
@@ 2.3. The period of Gushel–Mukai threefolds
Theorem 2.13 (Debarre–Kuznetsov).
@@
Two Gushel–Mukai threefolds are called period partner when they admit isomorphic double EPW sextics. Two Gushel–Mukai threefolds are called period dual when one associated double EPW sextics is the double dual EPW sextics for the other one.
Due to [7, Theorem 1.9], information of the period partners or duals for a Gushel–Mukai threefolds $X$ can be read from the Kuznetsov component ${\textbf{{Ku}}}(X)$. Once an addition data is included (serves as $V_{5}$), one can recover $X$ from ${\textbf{{Ku}}}(X)$ according to [29].

### v3

@@ Abstract.
We show that the double dual EPW sextic and double EPW sextic associated with a strongly smooth Gushel–Mukai surface can be realized as moduli spaces of semistable objects with respect to a stability condition on its bounded derived category. Also, we observe that the double dual EPW surface and double EPW surface associated with a special Gushel–Mukai threefold can be realized as moduli spaces of semistable objects on its Kuznetsov component. As an application, we refine a statement of Bayer and Perry about Gushel–Mukai threefolds with equivalent Kuznetsov components for the special ones.
@@
Gushel–Mukai varieties, moduli spaces, derived categories
@@
Primary 14D20 14E20 14J28 14J45 18G80
The first author is a member of the INdAM group GNSAGA (2025-2026). The second author is supported by SYSU Starting Grant No. 34000-12256019.
@@ 1. Introduction
Moduli spaces of semistable sheaves provide new methods to study the geometry of low dimension varieties and provide models for some higher dimensional varieties. Here we refer the readers to Huybrechts and Lehn [29] for a comprehensive account of this theory. On the other hand, moduli spaces of semistable sheaves on special members in a family of varieties are barely understood even for families of surfaces. Nevertheless, moduli spaces of semistable sheaves are in general only compatible with étale covers but not branched ones. These two disadvantages can be partly solved by considering instead the moduli spaces of semistable objects with respect to a notion of stability conditions introduced by Bridgeland [13], e.g. [37, 47].
@@ 1.1. GM varieties
An ordinary Gushel–Mukai surface is a smooth 2-dimensional transverse intersection $S=\operatorname{Gr}(2,V_{5})\cap\mathbb{P}(W)\cap Q$ where $V_{5}$ is a five-dimensional vector space, $W\subset\mathsf{\Lambda}^{2}V_{5}$ is a seven-dimensional vector subspace, and $Q\subset\mathbb{P}(W)$ is a quadratic hypersurface. Under the ample polarization $H$ given by $\mathscr{O}_{\mathbb{P}(W)}(1)|_{S}$, the surface $S$ is a Brill–Noether general K3 surface of degree $10$. Such a surface $S$ is called strongly smooth if the intersection $M_{S}:=\operatorname{Gr}(2,V_{5})\cap\mathbb{P}(W)$ is smooth. The double cover $X\rightarrow M_{S}$ branched over $S$ is called the special Gushel–Mukai threefold associated with $S$. Such a threefold $X$ has Picard rank $1$ and is Fano of index one and degree $10$ under the polarization $\mathscr{O}_{X}(1)$ given by pulling back $\mathscr{O}_{\mathbb{P}(W)}(1)$.
@@
of $X$, where $\mathscr{U}_{X}$ is the pullback of the tautological bundle $\mathscr{U}_{G}$ on $\operatorname{Gr}(2,V_{5})$. The associated equivariant category is equivalent to ${\textbf{{D}}}^{b}(S)$ according to [33, Section 8.2]. The dual cyclic group $\mathfrak{S}^{\vee}_{2}$ acts on ${\textbf{{D}}}^{b}(S)$ by an involutive autoequivalence $\Pi$ on ${\textbf{{D}}}^{b}(S)$ and the associated equivariant category is reversely equivalent to ${\textbf{{Ku}}}(X)$ by [22, Theorem 1.3].
According to [37, Section 6.1], one can find a stability condition $\sigma_{S}$ on ${\textbf{{D}}}^{b}(S)$ for any strongly smooth ordinary Gushel–Mukai surface $S$ such that the moduli spaces of semistable objects with respect to $\sigma_{S}$ detect many geometric features of $S$. Moreover, the stability condition $\sigma_{S}$ is compatible with the $\mathfrak{S}^{\vee}_{2}$-group action on ${\textbf{{D}}}^{b}(S)$ so that it induces a stability condition $\sigma_{X}$ on the equivariant category ${\textbf{{Ku}}}(X)$ and vice versa by virtue of [47, Section 4.3].
The forgetful functors between ${\textbf{{D}}}^{b}(S)$ and ${\textbf{{Ku}}}(X)$ induce covering morphisms between moduli spaces of semistable objects with respect to $\sigma_{S}$ and $\sigma_{X}$. The properties and first examples of these covering morphisms have been investigated in [37, 47]. In particular, some classic branched double covers can be reconstructed. This article aims to recover some EPW varieties associated with $S$ and $X$ as moduli spaces, expanding results in [37, Section 6].
@@ 1.2. EPW varieties
The vector space $V_{6}:=H^{0}(\mathbb{P}(W),\mathscr{I}_{S}(2))$ of quadrics in $W$ containing an ordinary Gushel–Mukai surface $S$ is six-dimensional and each ordinary Gushel–Mukai surface $S$ is canonically associated with a Lagrangian subspace $A(S)\subset\mathsf{\Lambda}^{3}V_{6}$ by [16]. One can construct a four-dimensional projective variety $\tilde{Y}^{\geq 1}_{A(S)}$ from $A(S)$, called the double EPW sextic associated with the Gushel–Mukai surface $S$ as in [41, 43]. The annihilator $A(S)^{\perp}\subset\mathsf{\Lambda}^{3}V^{\vee}_{6}$ of the Lagrangian subspace $A(S)$ is also a Lagrangian subspace and gives another four-dimensional projective variety $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ called the double dual EPW sextic [42] associated with $S$.
The singular locus of $\tilde{Y}^{\geq 1}_{A(S)}$ is an integral normal Cohen–Macaulay surface $Y^{\geq 2}_{A(S)}$ called the EPW surface associated with $S$. There exists a double cover over $Y^{\geq 2}_{A(S)}$ branched on its singular locus, which is called the double EPW surface and denoted by $\tilde{Y}^{\geq 2}_{A(S)}$ due to [20]. Similarly, one has the associated dual EPW surface $Y^{\geq 2}_{A(S)^{\perp}}$ and associated double dual EPW surface $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$.
@@ 1.3. Main results
In [37], the double sextics $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ and $\tilde{Y}^{\geq 1}_{A(S)}$ are realized as moduli spaces of semistable objects with respect to $\sigma_{S}$ for a general ordinary Gushel–Mukai surface $S$ based on the descriptions in [41, 44]. Here we provide the modular description of $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ for any strongly smooth ordinary Gushel–Mukai surface $S$ following [21, Section 7.5].
@@
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{1})$ of semistable objects with Mukai vector $v_{1}=(1,0,-1)$ is isomorphic to the double dual EPW sextic $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ associated with $S$.
However, an explicit birational model for $\tilde{Y}^{\geq 1}_{A(S)}$ is not available. In this case, one can pass to a strongly smooth ordinary Gushel–Mukai surface $S^{\prime}$ which is derived equivalent to the surface $S$ under a $\mathfrak{S}_{2}^{\vee}$-equivariant equivalence and satisfies $\tilde{Y}^{\geq 1}_{A(S)}\cong\tilde{Y}^{\geq 1}_{A(S^{\prime})^{\perp}}$ using [9, 44]. Then one can deduce the following statement using the results in [37, Section 6].
@@
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{2})$ of semistable objects with Mukai vector $v_{2}=(2,-H,2)$ is isomorphic to the double EPW sextic $\tilde{Y}^{\geq 1}_{A(S)}$ associated with $S$.
Moreover, the involutions on $\textsf{{M}}_{\sigma_{S}}(v_{1})$ and $\textsf{{M}}_{\sigma_{S}}(v_{2})$ induced by the $\mathfrak{S}_{2}^{\vee}$ group action coincide with the covering involutions on the corresponding EPW varieties. So according to [47], the inflation functor ${\textbf{{Ku}}}(X)\rightarrow{\textbf{{D}}}^{b}(S)$ induces two double covers
@@
The classes $\kappa_{1}$ and $\kappa_{2}$ form a basis for $\textsf{{N}}({\textbf{{Ku}}}(X))$ which is a rank two lattice such that an equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ either preserves or exchanges them up to a sign. It is conjectured in [34] and proved in [9, Theorem 5.12] using homological projective duality that two special Gushel–Mukai threefolds $X$ and $X^{\prime}$ have equivalent Kuznetsov components if and only if one has $\tilde{Y}^{\geq 1}_{A(X)}\cong\tilde{Y}^{\geq 1}_{A(X^{\prime})}$ or $\tilde{Y}^{\geq 1}_{A(X)}\cong\tilde{Y}^{\geq 1}_{A(X^{\prime})^{\perp}}$. Using the computations in [9, 37], a refinement can be established as an application of Theorem 1.1 and Proposition 1.2.
@@
Consider two special Gushel–Mukai threefolds $X$ and $X^{\prime}$, then $\tilde{Y}^{\geq 1}_{A(X)}\cong\tilde{Y}^{\geq 1}_{A(X^{\prime})}$ if and only if one has ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ sending $\kappa_{1}$ to $\kappa^{\prime}_{1}$ and $\kappa_{2}$ to $\kappa^{\prime}_{2}$ up to a sign; $\tilde{Y}^{\geq 1}_{A(X)}\cong\tilde{Y}^{\geq 1}_{A(X^{\prime})^{\perp}}$ if and only if one has ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ sending $\kappa_{1}$ to $\kappa^{\prime}_{2}$ and $\kappa_{2}$ to $\kappa^{\prime}_{1}$ up to a sign.
@@
At the end of this article, we exploit our results together with the Brill–Noether locus identification of a special Gushel–Mukai threefold in [32] to investigate the special Gushel–Mukai threefolds sharing the same Lagrangian subspaces.
@@ A Gushel–Mukai variety $X$ is a smooth $n$-dimensional intersection
Proposition 2.4 ([23, Lemma 2.7, Lemma 2.8]).
@@
In fact, due to [11, Remark 2.19], a Gushel–Mukai surface is possibly strongly smooth only when it is ordinary in the following definition.
@@ 2.2. The Lagrangian subspaces and EPW varieties
is canonically associated with a Lagrangian subspace $A(X)\subset\mathsf{\Lambda}^{3}V_{6}$ for $V_{6}=H^{0}(\mathbb{P}(W),\mathscr{I}_{X}(2))$ according to [18, Theorem 3.10]. One can define a series of projective varieties associated with the Lagrangian subspace. These varieties are introduced and intensively studied by O’Grady.
@@
can be endowed with a scheme structure $Y_{A}^{\geq\ell}$ for any integral $\ell$ according to [41, Section 2].
@@
A non-zero element of a Lagrangian subspace $A\subset\mathsf{\Lambda}^{3}V_{6}$ is called decomposable once it can be written in the form $v_{1}\land v_{2}\land v_{3}$.
@@
    the singular locus of $\tilde{Y}_{A}^{\geq 1}$ is the finite set $f^{-1}_{A}(Y_{A}^{\geq 3})$;
@@
    the variety $\tilde{Y}_{A}^{\geq 1}$ is a $K3^{[2]}$-type hyperkähler manifold once $Y_{A}^{\geq 3}=\varnothing$.
@@
$f_{A*}\mathscr{O}_{\tilde{Y}_{A}^{\geq 2}}\cong\mathscr{O}_{\tilde{Y}_{A}^{\geq 2}}\cong\omega_{Y_{A}^{\geq 2}}(-3)$
where $\mathscr{O}_{Y_{A}^{\geq 2}}(1)$ is induced by $\mathscr{O}_{\mathbb{P}(V_{6})}(1)$. Moreover, one has
@@
    the variety $\tilde{Y}_{A}^{\geq 2}$ is integral and normal;
@@
    the singular locus of $\tilde{Y}_{A}^{\geq 2}$ is the finite set $f^{-1}_{A}(Y_{A}^{\geq 3})$;
  • (d)
    the variety $\tilde{Y}_{A}^{\geq 2}$ is a smooth surface of general type once $Y_{A}^{\geq 3}=\varnothing$.
@@ 2.3. The period of Gushel–Mukai threefolds
Theorem 2.13 (Debarre–Kuznetsov [16]).
@@
Two Gushel–Mukai threefolds are called period partner when they admit isomorphic double EPW sextics. Two Gushel–Mukai threefolds are called period dual when one associated double EPW sextic is the double dual EPW sextic for the other one.
Due to [9, Theorem 1.9], information of the period partners or duals for a Gushel–Mukai threefold $X$ can be read from the Kuznetsov component ${\textbf{{Ku}}}(X)$. Once an additional data is included (serves as $V_{5}$), one can recover $X$ from ${\textbf{{Ku}}}(X)$ according to [32].